# Identifying Important Group of Pixels using Interactions

### v2

@@ 2 Related Work
Several recent studies have utilized a game-theoretic concept, interactions, to analyze various phenomena of deep learning models and quantify an effect of pixel cooperation on the model inference [7, 9, 21, 28, 29, 25]. Wang et al. [28] showed that the transferability of adversarial images has a negative correlation to the interactions. Zhang et al. [29] showed the similarity between the computation of interactions and dropout regularization. Deng et al. [9] discussed the difference in information obtained between humans and machine learning models using interactions. Sumiyasu et al. [25] investigated misclassification by models using interactions and discovered that the distribution of interactions varies with the type of misclassified images. Thus, interactions are helpful for understanding the model from the perspective of cooperative relationships between pixels. A critical issue of interaction-based analysis is its computational cost; the computation of interaction requires an exponential number of forward passes with respect to the number of pixels. In this paper, we propose an efficient approach to explain a model using variants of interactions (and also Shapley values), achieving the identification of important pixels with only a quadratic number of forward passes.
@@ 4.2 Pixel Deletion
$=\underset{b\in N\setminus\{b_{1}\}}{\mathrm{arg\,max}}\ \phi_{\mathrm{d}}^{(|N|)}(b)+(|N|-1)I_{\mathrm{d}}^{(|N|-1)}(b_{1},b).$ (12)
For $k\geq 3$, we can similarly show that minimizing $f(N\setminus S_{k-1}\cup\{b\})$ with respect to $b_{k}$ is equivalent to
$b_{k}$ $=\underset{b\in N\setminus S_{k-1}}{\operatorname{argmax}}\>[\phi_{\mathrm{d}}^{(|N|)}(b)$
$\quad+(|N|-|S_{k-1}|)\,I_{\mathrm{d}}^{(|N|-|S_{k-1}|)}(S_{k-1},b)].$ (13)
@@
Again, the greedy algorithm is described from a game-theoretic viewpoint. The only difference from the insertion case is that the interaction term is now weighted. Algorithm 2 summarises the procedure. The computational cost of the pixel deletion approach is the same as the pixel insertion approach, which only requires $\mathcal{O}(|N|^{2})$ times of forward passes in the worst case.
@@ 5 Experiments
Our experiments utilize the ImageNet dataset [10] and focus on analyzing Vision Transformer [11] pre-trained for the classification task. For baseline methods, we use Grad-CAM [23] [^2] The target layer of Grad-CAM is set to the one before the layer normalization in the final attention block of network. This choice is common, see https://github.com/jacobgil/pytorch-grad-cam., Grad-CAM++ [5], Attention rollout [1], Shapley values, and MoXI(-), which do not utilize the interactions present in MoXI. For insertion curve experiments, we use the Pixel Insertion approach, while for deletion curves, we utilize the Pixel Deletion approach. Following the previous studies [21, 29], we consider image patches instead of pixels to reduce computational costs. All methods calculate the contributions for $14\times 14$ patches with a patch size of $16\times 16$, which is equal to the patch size and the number of tokens in standard ViT models. We used a pre-trained ViT-T [^3] https://huggingface.co/WinKawaks/vit-tiny-patch16-224 [11], DeiT-T [^4] https://huggingface.co/facebook/deit-tiny-patch16-224 [26] and ResNet-18555https://huggingface.co/microsoft/resnet-18 [14]. We selected $1000$ images, one corresponding to each label, all of which were successfully classified in the test set. To reduce the computational burden, we computed Shapley values approximately by random sampling of $S$ in Eq. (1) as in other studies [4, 22, 29, 25]. The sampling size is set to 200. Moreover, we have adopted feature patch deletion as the masking method for Shapley values and interactions. In the following, we focus on ViT-T. See Appendix B for more results.
@@ G.1 Common corruptions
We implemented 19 types of common corruptions using the imagecorruptions module with severity 5.666 https://github.com/hendrycks/robustness. Figures 13 and 14 showcase the deletion curves with different corruptions for ViT-T and DeiT-T, respectively. The results demonstrate that our method gives a sharper decrease at the early stage of deletion curves than others, as in Sec. 5.3.

### v3

@@ 2 Related Work
Several recent studies have utilized a game-theoretic concept, interactions, to analyze various phenomena of deep learning models and quantify an effect of pixel cooperation on the model inference [9, 11, 23, 30, 31, 27]. [^1] There is also a similar approach to measure the second-order effects [4, 5], but this paper focuses on Shapley values and interactions, which are more widely used. Wang et al. [30] showed that the transferability of adversarial images has a negative correlation to the interactions. Zhang et al. [31] showed the similarity between the computation of interactions and dropout regularization. Deng et al. [11] discussed the difference in information obtained between humans and machine learning models using interactions. Sumiyasu et al. [27] investigated misclassification by models using interactions and discovered that the distribution of interactions varies with the type of misclassified images. Thus, interactions are helpful for understanding the model from the perspective of cooperative relationships between pixels. A critical issue of interaction-based analysis is its computational cost; the computation of interaction requires an exponential number of forward passes with respect to the number of pixels. In this paper, we propose an efficient approach to explain a model using variants of interactions (and also Shapley values), achieving the identification of important pixels with only a quadratic number of forward passes.
@@ 4.2 Pixel Deletion
$=\underset{b\in N\setminus\{b_{1}\}}{\mathrm{arg\,max}}\ \phi_{\mathrm{d}}^{(|N|-1)}(b\,|\,N\setminus\{b_{1}\})+\phi_{\mathrm{d}}^{(|N|-1)}(b_{1}\,|\,N\setminus\{b\})$
$\quad\quad\quad\quad\quad\quad+I_{\mathrm{d}}^{(|N|-1)}(b_{1},b)$ (12)
For $k\geq 3$, we can similarly show that minimizing $f(N\setminus S_{k-1}\cup\{b\})$ with respect to $b$ is equivalent to
$b_{k}$ $=\underset{b\in N\setminus S_{k-1}}{\operatorname{argmax}}\>[\phi_{\mathrm{d}}^{(|N|-1)}(b\,|\,N\setminus\{S_{k-1}\})$
$\quad\quad\quad\quad\quad\quad+\phi_{\mathrm{d}}^{(|N|-1)}(S_{k-1}\,|\,N\setminus\{b\})$
$\quad\quad\quad\quad\quad\quad+I_{\mathrm{d}}^{(|N|-|S_{k-1}|)}(S_{k-1},b)].$ (13)
@@
Again, the greedy algorithm is described from a game-theoretic viewpoint. In this case, we have two Shapley value terms. Algorithm 2 summarises the procedure. The computational cost of the pixel deletion approach is the same as the pixel insertion approach, which only requires $\mathcal{O}(|N|^{2})$ times of forward passes in the worst case.
@@ 5 Experiments
Our experiments utilize the ImageNet dataset [12] and focus on analyzing Vision Transformer [13] pre-trained for the classification task. For baseline methods, we use Grad-CAM [25] [^3] The target layer of Grad-CAM is set to the one before the layer normalization in the final attention block of network. This choice is common, see https://github.com/jacobgil/pytorch-grad-cam., Grad-CAM++ [7], Attention rollout [1], Shapley values, and MoXI(-), which do not utilize the interactions present in MoXI. For insertion curve experiments, we use the Pixel Insertion approach, while for deletion curves, we utilize the Pixel Deletion approach. Following the previous studies [23, 31], we consider image patches instead of pixels to reduce computational costs. All methods calculate the contributions for $14\times 14$ patches with a patch size of $16\times 16$, which is equal to the patch size and the number of tokens in standard ViT models. We used a pre-trained ViT-T [^4] https://huggingface.co/WinKawaks/vit-tiny-patch16-224 [13], DeiT-T [^5] https://huggingface.co/facebook/deit-tiny-patch16-224 [28] and ResNet-18666https://huggingface.co/microsoft/resnet-18 [16]. We selected $1000$ images, one corresponding to each label, all of which were successfully classified in the test set. To reduce the computational burden, we computed Shapley values approximately by random sampling of $S$ in Eq. (1) as in other studies [6, 24, 31, 27]. The sampling size is set to 200. Moreover, we have adopted feature patch deletion as the masking method for Shapley values and interactions. In the following, we focus on ViT-T. See Appendix B for more results.
@@ G.1 Common corruptions
We implemented 19 types of common corruptions using the imagecorruptions module with severity 5.777 https://github.com/hendrycks/robustness. Figures 13 and 14 showcase the deletion curves with different corruptions for ViT-T and DeiT-T, respectively. The results demonstrate that our method gives a sharper decrease at the early stage of deletion curves than others, as in Sec. 5.3.